# Can GenAI Move from Individual Use to Collaborative Work? Experiences, Challenges, and Opportunities of Integrating GenAI into Collaborative Newsroom Routines

### v2

@@ 6.1. Structural Challenges: Lack of Organizational Mechanisms for Integration
In addition to traceability issues, GenAI tools lack shared interfaces that support collaborative review and editing. Unlike tools such as Google Docs, which facilitate real-time co-editing and inline commenting, most GenAI platforms are designed for single-user interaction. Journalists often copy and paste GenAI outputs into shared documents or use them as prompts for team discussions without revealing the prompts they used, the iterations they tried, or any potential misunderstandings by the model. These limitations are compounded by the fact that GenAI is fundamentally a black-box system. Editors, especially senior editors without a technical background, often lack the AI literacy needed to critically evaluate GenAI outputs. As P3 remarked, “On one hand, we don’t know how the journalist interacted with the GenAI. On the other hand, we don’t know how the GenAI came up with its response. It’s like two black boxes stacked on top of each other, which creates real challenges for editors.” This dual opacity of both human prompting and machine output undermines editorial confidence and creates a major barrier to integrating GenAI into collaborative newsroom routines.
@@
The organizational structure of most newsroom tech teams further exacerbates these technical constraints. In our interviews, P1, P2, and P3 all acknowledged that their teams of developers were typically small, under-resourced, and primarily focused on building core GenAI functionalities, such as fine-tuning models or ensuring data compliance, rather than designing interfaces that support multi-user collaboration. As P1 explained, “Right now we’re still at the stage of putting food on the table, not worrying about how good the meal is. We need to get a basic GenAI platform running internally in the newsroom first to meet Chinese government’s expectations.” P1 further mentioned, “Only then can we think about building collaborative platforms. So for now, we let journalists explore freely and use whichever GenAI tools they prefer, instead of developing a system for coordinated use. We just don’t have the capacity.” This prioritization of backend capabilities over collaborative usability limits the newsroom’s ability to develop traceable and shared GenAI workflows.
@@
This enhances the challenges for front-line journalist-editors collaborations around GenAI and the hesitations to promote an official guideline to integrate GenAI into newswork collaborations. Editors are no longer only concerned with factual accuracy or stylistic quality; they must also anticipate how GenAI usage will be perceived by increasingly critical audiences.
@@
P14, for example, regularly monitors discussions on Chinese social media platforms such as Xiaohongshu and Weibo. She observed that readers are becoming more vigilant and suspicious about the presence of GenAI in journalistic content. As she described, “In the past, it was enough to ensure that the content met quality standards. Audiences wouldn’t notice or care whether AI was involved. But this year, especially in 2025, that has changed. You’ll find a lot of social media users actively trying to identify whether certain news articles were written by GenAI. And if they think the content was AI-generated, they feel disappointed. Sometimes these posts go viral, with thousands of likes, shares, and comments.”
This shift in public scrutiny introduces reputational risks for newsrooms and adds further pressure on editorial teams to detect GenAI usage, that are already structurally and technically difficult to manage, as discussed earlier. Without systematic mechanisms for tracing GenAI contributions or standard protocols for disclosure, editors face heightened audience expectations with limited organizational support.
As a result, newsroom leaders such as P1 and P2 are particularly cautious about openly integrating GenAI into editorial routines. Many fear that any premature or poorly safeguarded disclosure of GenAI use could trigger criticism, undermine credibility, and damage institutional trust. This apprehension leads organizations to delay or avoid formalizing GenAI practices, as external scrutiny amplifies internal hesitation and compounds the difficulty of embedding GenAI in collaborative workflows.
@@
6.2. Cultural Challenges: Rethinking Journalists’ Identities in GenAI-Augmented Collaborative Work
@@
Culturally, journalists often perceive the use of GenAI as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to journalist peers and editorial supervisors. This secrecy poses a major obstacle to integrating GenAI into collaborative newsroom workflows. Many journalists do not yet regard GenAI as a legitimate component of professional practice, but rather as a gray-area personal assistant. Because these practices are seen as outside the boundaries of “real journalism,” they are hidden from view, fragmenting knowledge and preventing the development of shared norms.
@@
P24 explicitly described using GenAI as a form of professional shame, likening it to “cutting corners when no one is watching.” She explained that even if the final output is polished through human editing, the mere fact that the first draft was machine-generated made her feel like she had “cheated on an exam.” For her, GenAI was not a neutral tool, but a shortcut that undermined the pride she took in crafting language herself. “It’s like copying someone else’s homework and rewriting it to make it sound like your own,” she added. “Even if no one finds out, you still know what you did.”
@@
Other metaphors further reinforced this sense of concealment. P27 referred to using GenAI as “ghostwriting.” She remarked, “Ghostwriting is not uncommon in journalism, especially in commercial collaborations. In many Chinese news reports involving corporate clients, the company drafts the content and the newspaper simply publishes it under a journalist’s byline. But it’s never openly discussed. It’s a hidden shame. If we don’t talk about ghostwriting, why would we publicly talk about GenAI?” In this framing, GenAI’s most powerful capability—text generation—is morally likened to unethical journalistic practices such as plagiarism or covert authorship and is not celebrated as innovation. These comparisons reflect not only individual discomfort, but also the normative environment in which journalists operate: one in which journalistic value is rooted in human creativity, intellectual labor, originality, and textual craftsmanship.
@@
Many participants expressed concern that GenAI could devalue the profession itself, transforming journalism from a human enterprise into something replaceable by machines. As P24 self-mockingly noted, “We often talk about what professions AI might replace. News reports usually say it will replace low-creativity jobs like data entry. But if even high-literacy, high-creativity journalism can be assisted by GenAI, doesn’t that say something about how little society values our work?” T2 shared a widely discussed case in the Chinese news industry that illustrates this anxiety. T2 said a prominent Chinese newsroom recently sought to develop its own GenAI system for news writing and asked journalists to contribute training data by submitting ten of their most accomplished past articles. Yet, almost no journalist complied. As T2 explained, journalists feared that handing over their best work would not only strip them of control over the copyright of their writing but also allow the system to absorb their style and voice, erasing the individuality of their labor. For many, the idea that their copyrighted articles could be repurposed as raw material for a machine epitomized the risk of being replaced by the very system trained on their work. This moral framing of authorship, as something earned through struggle and talent, helps explain why GenAI remains a taboo topic, even within news organizations that encourage experimentation.
@@
Fear of peer and supervisory judgment also played a critical role. In environments where editorial authority is tied to experience and skill, admitting to AI assistance might be interpreted as a lack of competence or diligence. Junior journalists especially felt that revealing GenAI usage might be seen as laziness or unprofessionalism. P27 reflected, “My editor is very old-school. If I told him I used AI to structure a piece or just organize interview notes, he’d think I wasn’t taking the job seriously.” This perception of stigma reinforces secrecy and prevents GenAI from being openly incorporated into collaborative work, even when it could enhance efficiency or creativity.
Additionally, the experimental nature of GenAI tools exacerbates their cultural marginalization. Many journalists are still figuring out how to use these tools effectively, and are therefore reluctant to share incomplete or uncertain practices. P25 expressed embarrassment about this phase of exploration: “I don’t know if other journalists are using GenAI. My computer skills are very low. I’m just playing around.” This hesitation further isolates GenAI use and prevents organizations from building collective intelligence around GenAI integration.
@@
7. Findings: Opportunities Toward Future GenAI-Augmented Newswork Collaboration (RQ3)
@@
Our participants further articulated concrete visions for how GenAI could be embedded into the future of collaborative newswork. Our findings show that journalists anticipate GenAI-augmented collaboration will require rethinking organizational priorities (subsection 7.1), fostering a culture of responsible use (subsection 7.2), and developing better infrastructures (subsection 7.3) and new routines (subsection 7.4) that embed GenAI into collaborative practice.
@@ 7.1. Rethinking Organizational Prioritization: Making Collaboration a Core Design Goal
As P1 explained, “When resources are limited and especially when the government wants to see concrete GenAI system design outcomes, such as a functional prototype system that can be showcased in public or to the government as evidence of progress.” This mindset, while understandable in a results-driven institutional context like China’s, implicitly relegates collaboration to a downstream concern: something to be added later, once the core AI infrastructure is in place, rather than a foundational element of system design.
To address this imbalance, participants called for a cultural shift in GenAI development: collaboration must be treated as a first-order concern, not an afterthought. They outlined three specific strategies for achieving this transformation: First, involving editors and journalists early in the development process. Rather than viewing users as passive recipients of prebuilt systems, participants argued that journalists must be embedded in the development cycle from the beginning to ensure GenAI tools are grounded in collaborative realities. As P18 put it, “If no one is thinking about how these tools work in collaboration, how our colleagues or editors might perceive their use, it’s hard to trust that the tools are reliable in a newsroom collaboration context.”
Second, embed collaboration into the technical design of GenAI systems. Participants emphasized that GenAI platforms must support collaborative features, such as shared editing environments, real-time commenting, and visible histories of GenAI interactions. As P8 noted, “As an editor, I need to know how journalists interact with GenAI. If that visibility were built into the platform, it would help. But without it, we obviously can’t monitor every journalist’s chat history with GenAI, that would be outrageous and a violation of privacy.”
Third, expand GenAI evaluation metrics to include collaboration outcomes. Currently, internal GenAI pilot programs and evaluations often focus narrowly on performance: how well the system generates news text, images, or video. But participants argued that such metrics miss a critical dimension: whether GenAI improves or undermines collaborative newswork. As P12 suggested, “I’m not sure how to measure it, but maybe future evaluations should also include whether these tools actually enhance or hurt teamwork.”
These reflections point to a broader organizational insight for future GenAI design: its true value lies in how well it supports collaborative work in journalism: editorial judgment, team coordination, and collective accountability.
@@ 7.2. Fostering a Cultural Shift in Newsrooms: Embracing Responsible GenAI Use as Legitimate Newswork
Several participants called for a broader cultural transformation within their organizations: the normalization and legitimization of responsible GenAI use in routine journalistic practices. Although GenAI tools are already widely adopted by front-line journalists and editors individually, participants noted that their use remains largely hidden, ad hoc, or even stigmatized. Journalists worry that acknowledging reliance on GenAI might be interpreted as laziness, a lack of editorial rigor, or diminished professional credibility, especially in legacy newsrooms where traditional reporting values continue to dominate.
To address this hesitation, participants emphasized the importance of building a newsroom culture in which responsible GenAI use is not only accepted but also recognized as a legitimate form of editorial labor. P20 remarked, “We shouldn’t pretend everyone is doing everything by hand. That’s not honest, and it discourages people from trying. The reality is, many journalists are already using GenAI. We can’t pretend no one is.” P20 further explained that if GenAI use is becoming inevitable, then the next step must be to promote responsible use, not prohibit it, particularly given the practical impossibility of policing private GenAI use on personal devices.
To support this shift, participants urged newsroom leadership to provide clear internal messaging: the use of AI tools, when transparent, accountable, and thoughtfully integrated, should not be seen as undermining journalistic professionalism. In contrast, such use may reflect a strategic and informed approach to emerging technologies in pursuit of journalistic goals.
Participants also suggested that this cultural shift requires redefining what it means to be a responsible journalist in the GenAI era. Instead of viewing journalistic integrity as the absence of AI involvement, integrity should be evaluated based on the transparency and oversight exercised in using these tools. P3 emphasized, “Responsible GenAI use should be part of our professional standards, not something we hide from our peers or our audience.”
In participants’ vision, this transformation would also create space for open dialogue and peer learning around GenAI practices. “Journalists could more freely exchange prompt techniques, usage tips, and editorial guidelines” (P8), while “teams could collaboratively reflect on how GenAI use aligns with shared values” (P12).
Ultimately, legitimizing responsible GenAI use is a necessary step toward building a more adaptive, collaborative, and future-oriented newsroom culture.
@@ 7.3. Embedding GenAI into Newsroom Infrastructures for Future Collaboration
In contrast, according to our participants, embedding GenAI functionalities directly into newsrooms’ internal content management systems could potentially help transform GenAI from a siloed tool into an auditable and collectively accountable part of journalistic work. According to P3, if all interactions, such as prompts, generated drafts and post-editing revisions, were automatically captured and stored on a central platform of the newsroom, editors would be able to trace AI involvement, clarify accountability and make informed editorial decisions in the future. P3 put it: “We already have plans to build our own in-house GenAI system. What we need now is simply to integrate how people are already using GenAI into that system.” P1 agreed that, news organizations already investing in GenAI development can extend that investment to support collaborative visibility and editorial control. Such an integration also safeguards collaborative routines. It enables editorial teams to work from a common source of truth, reduces the burden on journalists to self-report AI use, and mitigates the reputational risks of AI-generated content slipping through unnoticed.
@@
In this context, the integration of GenAI is also a way to maintain institutional legitimacy and prepare for potential government audits or public scrutiny regarding how AI is used in sensitive editorial decisions. In some state-run newsrooms, such as CCTV and CGTN, every piece of news must already pass through a centralized editorial platform that records each publication decision. As P1 explained, “We already have content management systems where every news article is tracked and every editorial decision is logged. So it makes sense to extend that to AI as well, especially in China, where we need to be prepared to justify our editorial decisions at any time, whether to the government or to the public.”
@@
In P1’ imaginations, this kind of integration supports risk mitigation at scale. By embedding GenAI into existing editorial infrastructures, media organizations can respond more quickly to reputational, legal, or political concerns if AI-generated content is later questioned. It also creates a shared point of accountability for team members, strengthening collective responsibility and reducing the likelihood that GenAI use remains a hidden, individualized practice outside the editorial chain of command.
7.4. Creating New Work Routines around GenAI

### v3

@@ 6.1. Structural Challenges: Lack of Organizational Mechanisms for Integration
In addition to traceability issues, GenAI tools lack shared interfaces that support collaborative review and editing. Unlike tools such as Google Docs, which facilitate real-time co-editing and inline commenting, most GenAI platforms are designed for single-user interaction. Journalists often copy and paste GenAI outputs into shared documents or use them as prompts for team discussions without revealing the prompts they used, the iterations they tried, or any potential misunderstandings by the model. These limitations are compounded by the fact that GenAI is fundamentally a black-box system. Editors, especially senior editors without a technical background, often lack the AI literacy needed to critically evaluate GenAI outputs. As $P3_{\text{male, editor/manager}}$ remarked, “On one hand, we don’t know how the journalist interacted with the GenAI. On the other hand, we don’t know how the GenAI came up with its response. It’s like two black boxes stacked on top of each other, which creates real challenges for editors.” This “double black box”—the opacity of human interaction layered atop the opacity of model reasoning—captures a unique form, where neither the human process nor the machine process is inspectable, auditable, or mutually accountable. This dual opacity of both human prompting and machine output undermines editorial confidence and creates a major barrier to coordinate GenAI into collaborative newswork.
@@
The organizational structure of most newsroom tech teams further exacerbates these technical constraints. In our interviews, $P1_{\text{female, editor/manager}}$, $P2_{\text{female, editor/manager}}$, and $P3_{\text{male, editor/manager}}$ all acknowledged that their teams of developers were typically small, under-resourced, and primarily focused on building core GenAI functionalities, such as fine-tuning models or ensuring data compliance, rather than designing interfaces that support multi-user collaboration. As $P1_{\text{female, editor/manager}}$ explained, “Right now we’re still at the stage of putting food on the table, not worrying about how good the meal is. We need to get a basic GenAI platform running internally in the newsroom first to meet Chinese government’s expectations.” $P1_{\text{female, editor/manager}}$ further mentioned, “Only then can we think about building collaborative platforms. So for now, we let journalists explore freely and use whichever GenAI tools they prefer, instead of developing a system for coordinated use. We just don’t have the capacity.”
@@
This enhances the challenges for front-line journalist-editors collaborations around GenAI and the hesitations to promote an official guideline to coordinate GenAI into collaborative newswork. Editors are no longer only concerned with factual accuracy or stylistic quality; they could also anticipate how GenAI usage will be perceived by increasingly critical audiences.
@@
$P14_{\text{female, editor/journalist}}$, for example, regularly monitors discussions on Chinese social media platforms such as Xiaohongshu and Weibo. She observed that readers are becoming more vigilant and suspicious about the presence of GenAI in journalistic content. As she described, “In the past, it was enough to ensure that the content met quality standards. Audiences wouldn’t notice or care whether AI was involved. But this year, especially in 2025, that has changed. You’ll find a lot of social media users actively trying to identify whether certain news articles were written by GenAI. And if they think the content was AI-generated, they feel disappointed. Sometimes these posts go viral, with thousands of likes, shares, and comments.”
This shift in public scrutiny introduces reputational risks for newsrooms and adds further pressure on editorial teams to detect GenAI usage, that are already structurally and technically difficult to manage, as discussed earlier. As a result, newsroom leaders such as $P1_{\text{female, editor/manager}}$ and $P2_{\text{female, editor/manager}}$ are particularly cautious about openly coordinating GenAI into editorial newswork. Many fear that any premature or poorly safeguarded disclosure of GenAI use could trigger criticism, undermine credibility, and damage institutional trust.
@@
6.2. Cultural Challenges: Rethinking Journalists’ Identities in Collaborative Newswork Around GenAI
@@
Culturally, journalists often perceive the use of GenAI as private, experimental, or even stigmatized, leading to reluctance in disclosing their practices or outputs to journalist peers and editorial supervisors. This secrecy poses a major obstacle to coordinating GenAI into collaborative newsroom workflows. Many journalists do not yet regard GenAI as a legitimate component of professional practice, but rather as a gray-area personal assistant. Because these practices are seen as outside the boundaries of “real journalism,” they are hidden from view, preventing the development of shared norms.
@@
$P24_{\text{female, journalist}}$ explicitly described using GenAI as a form of professional shame, likening it to “cutting corners when no one is watching.” She explained that even if the final output is polished through human editing, the mere fact that the first draft was machine-generated made her feel like she had “cheated on an exam.” For her, GenAI was not a neutral tool, but a shortcut that undermined the pride she took in crafting language herself. “It’s like copying someone else’s homework and rewriting it to make it sound like your own,” she added. “Even if no one finds out, you still know what you did.”
@@
Other metaphors further reinforced this sense of concealment. $P27_{\text{male, journalist}}$ referred to using GenAI as “ghostwriting.” She remarked, “Ghostwriting is not uncommon in journalism, especially in commercial collaborations. In many Chinese news reports involving corporate clients, the company drafts the content and the newspaper simply publishes it under a journalist’s byline. But it’s never openly discussed. It’s a hidden shame. If we don’t talk about ghostwriting, why would we publicly talk about GenAI?” In this framing, GenAI’s most powerful capability—text generation—is morally likened to unethical journalistic practices such as plagiarism or covert authorship and is not celebrated as innovation. These comparisons reflect not only individual discomfort, but also the normative environment in which journalists operate: one in which journalistic value is rooted in human creativity, intellectual labor, originality, and textual craftsmanship.
Many participants expressed concern that GenAI could devalue the profession itself, transforming journalism from a human enterprise into something replaceable by machines. As $P24_{\text{female, journalist}}$ self-mockingly noted, “We often talk about what professions AI might replace. News reports usually say it will replace low-creativity jobs like data entry. But if even high-literacy, high-creativity journalism can be assisted by GenAI, doesn’t that say something about how little society values our work?”
@@
$P3_{\text{male, editor/manager}}$ shared a widely discussed case in the Chinese news industry that illustrates this anxiety. $P3_{\text{male, editor/manager}}$ said a prominent Chinese newsroom recently sought to develop its own GenAI system for news writing and asked journalists to contribute training data by submitting ten of their most accomplished past articles. Yet, almost no journalist complied. As $P3_{\text{male, editor/manager}}$ explained, journalists feared that handing over their best work would not only strip them of control over the copyright of their writing but also allow the system to absorb their style and voice, erasing the individuality of their labor. For many, the idea that their copyrighted articles could be repurposed as raw material for a machine epitomized the risk of being replaced by the very system trained on their work. This moral framing of authorship, as something earned through struggle and talent, helps explain why GenAI remains a taboo topic, even within news organizations that encourage experimentation.
@@
Fear of peer and supervisory judgment also played a critical role. In environments where editorial authority is tied to experience and skill, admitting to AI assistance might be interpreted as a lack of competence or diligence. Junior journalists especially felt that revealing GenAI usage might be seen as laziness or unprofessionalism. $P27_{\text{male, journalist}}$ reflected, “My editor is very old-school. If I told him I used AI to structure a piece or just organize interview notes, he’d think I wasn’t taking the job seriously.” This perception of stigma reinforces secrecy and prevents GenAI from being openly coordinated into collaborative work, even when it could enhance efficiency or creativity.
@@
7. Findings: Opportunities Toward Future Coordination of GenAI into Collaborative Newswork (RQ3)
@@
Our participants further articulated concrete visions for how GenAI could be embedded into the future of collaborative newswork. Our findings show that journalists anticipate coordinating GenAI in future collaboration will require rethinking organizational priorities (subsection 7.1), fostering a culture of responsible use (subsection 7.2), and developing better infrastructures (subsection 7.3) and new workflow (subsection 7.4) that embed GenAI into collaborative practice.
@@ 7.1. Rethinking Organizational Prioritization: Making Collaboration a Core Design Goal
As $P1_{\text{female, editor/manager}}$, who oversees organizational-level technology and workflow decisions in her newsroom, explained, “When resources are limited and especially when the government wants to see concrete GenAI system design outcomes, such as a functional prototype system that can be showcased in public or to the government as evidence of progress.” This mindset, while understandable in a results-driven institutional context like China’s, implicitly relegates collaboration to a downstream concern: something to be added later, once the core AI infrastructure is in place, rather than a foundational element of system design.
To address this imbalance, participants called for a cultural shift in GenAI development: collaboration could be treated as a first-order concern, not an afterthought. They outlined three specific strategies for achieving this transformation: First, involving editors and journalists early in the development process. Rather than viewing users as passive recipients of prebuilt systems, participants argued that journalists could be embedded in the development cycle from the beginning to ensure GenAI tools are grounded in collaborative realities. As $P18_{\text{male, editor/journalist}}$ put it, “If no one is thinking about how these tools work in collaboration, how our colleagues or editors might perceive their use, it’s hard to trust that the tools are reliable in a newsroom collaboration context.”
Second, embed collaboration into the technical design of GenAI systems. Participants emphasized that GenAI platforms could support collaborative features, such as shared editing environments, real-time commenting, and visible histories of GenAI interactions. As $P8_{\text{male, editor}}$ noted, “As an editor, I need to know how journalists interact with GenAI. If that visibility were built into the platform, it would help. But without it, we obviously can’t monitor every journalist’s chat history with their GenAI, that would be outrageous and a violation of privacy.”
Third, expand GenAI evaluation metrics to include collaboration outcomes. Currently, internal GenAI pilot programs and evaluations often focus narrowly on performance: how well the system generates news text, images, or video. But participants argued that such metrics miss a critical dimension: whether GenAI improves or undermines collaborative newswork. As $P12_{\text{female, journalist}}$ suggested, “I’m not sure how to measure it, but maybe future evaluations could also include whether these tools actually enhance or hurt teamwork.”
@@ 7.2. Fostering a Cultural Shift in Newsrooms: Embracing Responsible GenAI Use as Legitimate Newswork
Several participants called for a broader cultural transformation within their organizations: the normalization and legitimization of responsible GenAI use in journalistic practices. Although GenAI tools are already widely adopted by front-line journalists and editors individually, participants noted that their use remains largely hidden, ad hoc, or even stigmatized. Journalists worry that acknowledging reliance on GenAI might be interpreted as laziness, a lack of editorial rigor, or diminished professional credibility, especially in legacy newsrooms where traditional reporting values continue to dominate.
To address this hesitation, participants emphasized the importance of building a newsroom culture in which responsible GenAI use is not only accepted but also recognized as a legitimate form of editorial labor. $P20_{\text{male, journalist}}$ remarked, “We couldn’t pretend everyone is doing everything by hand. That’s not honest, and it discourages people from trying. The reality is, many journalists are already using GenAI. We can’t pretend no one is.” $P20_{\text{male, journalist}}$ further explained that if GenAI use is becoming inevitable, then the next step could be to promote responsible use, not prohibit it, particularly given the practical impossibility of policing private GenAI use on personal devices.
To support this shift, participants urged newsroom leadership to provide clear internal messaging: the use of AI tools, when transparent, accountable, and thoughtfully coordinated, could not be seen as undermining journalistic professionalism. In contrast, such use may reflect a strategic and informed approach to emerging technologies in pursuit of journalistic goals.
Participants also suggested that this cultural shift requires redefining what it means to be a responsible journalist in the GenAI era. Instead of viewing journalistic integrity as the absence of AI involvement, integrity could be evaluated based on the transparency and oversight exercised in using these tools. $P3_{\text{male, editor/manager}}$ emphasized, “Responsible GenAI use could be part of our professional standards, not something we hide from our peers or our audience.”
In participants’ vision, this transformation would also create space for open dialogue and peer learning around GenAI practices. “Journalists could more freely exchange prompt techniques, usage tips, and editorial guidelines” ($P8_{\text{male, editor}}$), while “teams could collaboratively reflect on how GenAI use aligns with shared values” ($P12_{\text{female, journalist}}$).
@@ 7.3. Embedding GenAI into Newsroom Infrastructures for Future Collaboration
In contrast, according to our participants, embedding GenAI functionalities directly into newsrooms’ internal content management systems could potentially help transform GenAI from a siloed tool into an auditable and collectively accountable part of journalistic work. According to $P3_{\text{male, editor/manager}}$, if all interactions, such as prompts, generated drafts and post-editing revisions, were automatically captured and stored on a central platform of the newsroom, editors would be able to trace AI involvement, clarify accountability and make informed editorial decisions in the future. $P1_{\text{female, editor/manager}}$ agreed that, news organizations already investing in GenAI development can extend that investment to support collaborative visibility and editorial control.
@@
In this context, the coordination of GenAI is also a way to maintain institutional legitimacy and prepare for potential government audits or public scrutiny regarding how AI is used in sensitive editorial decisions. In some state-run newsrooms, such as CCTV and CGTN, every piece of news could already pass through a centralized editorial platform that records each publication decision (Shen and Zhou, 2025; China Computer Federation, 2024). As $P1_{\text{female, editor/manager}}$ explained, “We already have content management systems where every news article is tracked and every editorial decision is logged. So it makes sense to extend that to AI as well, especially in China, where we need to be prepared to justify our editorial decisions at any time, whether to the government or to the public.”
@@
In $P1_{\text{female, editor/manager}}$’s imaginations, this kind of coordination supports risk mitigation at scale. By embedding GenAI into existing editorial infrastructures, media organizations can respond more quickly to reputational, legal, or political concerns if AI-generated content is later questioned. It also creates a shared point of accountability for team members, strengthening collective responsibility and reducing the likelihood that GenAI use remains a hidden, individualized practice outside the editorial chain of command.
7.4. Creating New Workflow Around GenAI